OpenStudy (anonymous):

graph f(x)= 2x+5 if x <0 5 if 0 ≤ x ≤ 10 2x-10 if x > 10

OpenStudy (freckles):

let's do this in pieces We need to graph the line y=2x+5 for x<0

OpenStudy (freckles):

So this line we will only draw on the x interval (-inf,0) now this piece doesn't exist at x=0 but it does include all values to the left of x=0 so I'm still going to use x=0 to find where I should stop the line for y-coordinate and then use whatever value to the left of x=0 to finish drawing the line

OpenStudy (freckles):

so y=2(0)+5=? now also choosing same value to the left of x=0... y=2(-2)+5=?

OpenStudy (anonymous):

Okay so then y=2(-2)+5 y= 1?

OpenStudy (freckles):